OpenStudy (anonymous):

**CALC II** Find the volume of the solid generated by revolving the described region about the given axis: The region in the first quadrant enclosed on the left by the y-axis and on the right by the curve x = 4 (y^2 - y^3), rotated about the (a)x-axis; (b)y = 1; (c) y = 2; (d) y = -2

OpenStudy (nincompoop):

show us your ATTEMPT

OpenStudy (nincompoop):

gosh, you're doing calculus now, act more mature about doing work

OpenStudy (aum):

x = 4 (y^2 - y^3) = 4y^2(1 - y) Find x-intercept(s): Set y = 0. x = 0 Find x-intercepts: Set x = 0. 0 = 4y^2(1 - y). y = 0, y = 1. Next, a rough sketch:

OpenStudy (aum):